OpenStudy (anonymous):

lilg132 can you please help me on my algebra again???

OpenStudy (anonymous):

ok but i think you are a good tutor

OpenStudy (anonymous):

{3x-2y=7 {x+3y=-5

OpenStudy (anonymous):

Are you still there?

OpenStudy (lilg132):

ok so lets use the second equation for substitution and thanks x = -5 - 3y now lets use this equation with the first equation and replace and x's with that so 3x - 2y = 7 will now be 3(-5 - 3y) - 2y = 7 which is -15 - 9y - 2y = 7 bring the numbers to one side and the variables to the other (remember bringing a number to the other side changes its value e.g. bringint -15 to the other side will change it to +15 ) -9y - 2y = 7 + 15 -11y = 22 now to get y on its own we divide it by the number next to y which is 11 and you also divide the other side by the same number so -11y / 11 = 22 / 11 -y = 2 to get -y into a positive we just flip it around (remember what i said bringin anything to the other side will change its value e.g. -y to the other side will change it to +y or y) -2 = y that is the answer to y now can you use that to work out x? if not please tell me and i will show you the steps again

OpenStudy (anonymous):

i need the x i dont get it

OpenStudy (lilg132):

ok i will help you one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh we got the x sorry

OpenStudy (anonymous):

ok if you help me on one more question i swear it will be the last one

OpenStudy (lilg132):

now we have the answer to y which is y = -2 now pick any of the equations and where ever you see y replace it with -2 then solve the equation i will do this for you this time and then try doing the next question on your own if you understand if you get stuck do tell me. lets use the first equation 3x-2y=7 remember replace the y's with tha answer we got which was -2 so 3x - 2(-2) = 7 3x + 4 = 7 (remember - * - = +) now we bring the numbers to one side and have the variables on the other (dont forget bringing a number to the other side will change its value e.g. bringing +4 to the other side will change it to -4) 3x = 7 - 4 3x = 3 now we divide both sides by the number next to x to get x on its own and to solve x so 3x / 3 = 3 / 3 x = 1 do you understand this ?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

that is why we are asking you for help

OpenStudy (lilg132):

ok can you tell me which part you dont understand

OpenStudy (anonymous):

i dont understand how to figure out y

OpenStudy (anonymous):

i just need you to help me on one more question but to only help me on y because i already know how to figure out x

OpenStudy (anonymous):

the question is {-2x+y=0 {5x+3y=-11

OpenStudy (lilg132):

ok so lets use the first equation because there is nothing next to y so its easier now the first equation is -2x + y = 0 we rearrange this to get y = but to get y on its own we bring everything else to the other side so y = 0 + 2x that is how you get y now the word substitute is another way of saying instead of so everytime you see Y instead of keeping y in the equation replace it with what we have y = 0 + 2x so lets use the second equation which is 5x+3y=-11 now we see y in this equation but we replace it by what we found before so 5x + 3(0 + 2x) = -11 now you expand that bracket first 3 * 0 = 0 3 * 2x = 6x so the equation should now look like this 5x + 0 + 6x = -11 we can take out the 0 as it wont affect the answers in any way so we have 5x + 6x = -11 now we just add the two variables up which gives you 11x = -11 now remember to get x on its own and to solve it you need to divide both sides by the number next to x which is 11 so 11x / 11 = -11 / 11 is x = -1

OpenStudy (anonymous):

ok thanks so much, we appreciate your amazing tutoring

OpenStudy (lilg132):

now can you use this to work out y so now pick any of the two equations in your question and everywhere you see x replace it with the value we got which is -1

OpenStudy (lilg132):

your welcome